# Neural Kinematic Bases for Fluids

### v1

@@ 4.3 Ablation study
Two-Stage
Figure 11: Bases generated by omitting different parts of our pipeline. Each omission leads to different artifacts.
@@
$t=1.8s$
$t=2.9s$
$t=4.7s$
Figure 12: Example of different fluid animation frames for different input sketches. The flow adapts to the position of the circles and the curves and produces natural animation.
We showcase the importance of three main features of our pipeline (Figure 11). To ease the comparison, we limit the dataset to 3 different circles as we do not test for generalization.
Smoothness loss
This is the only loss that is not motivated by fundamental physical law. We included it to bias the MLP towards smooth solutions (which our bases must be). Adding it improves the reliability of the training process: in our experiments, without it, we obtain unreasonable bases more frequently (around 30% more).
Rounded Corners
Two-stage training for boundary
As previously mentioned, we train our MLP network in two phases. Without this approach, the bases do not align with the boundary (for power of 2) or lack smoothness (for power of 4).
4.4 Animations
We run all our experiments on a GeForce RTX 3090. Generating the initial fluid is interactive, while the simulation runs at 40 frames per second for 250 thousand points.
Two-dimensional
@@ 4.4 Animations
All our methods can be generalized to three dimensions. The domain is now represented by three spheres and their radii (12 parameters instead of 9); the MLP network accepts 3D points and produces a volumetric vector field. Since all losses are physically motivated, they naturally extend to arbitrary dimensions. Similarly, the advection of the fluid is dimension agnostic. Since sketching volumetric curves is more complicated, we initialize the flow using random $\alpha$. Figure 13 shows the results for three overlapping spheres, while Figure 1 shows how the flow passes between the spherical holes. Both animations are run in real-time as the MLP is only evaluated. Both figures show the velocity field both as arrows and streamlines colored by their velocity.
@@
$t=10s$
Figure 13: Three-dimensional fluid animation using our neural kinematic bases. We show both the velocity and streamlines colored by their velocity.
5 Conclusions
We introduce a novel neural kinematic base represented by an MLP that captures fundamental physical properties and generalizes to different domains. We show how using our bases and a standard advection integrator, we can generate two- and three-dimensional animations.
To keep the training times reasonable, we decided to use only three circles and ten bases, but our method should be able to generate more bases for more complex domains. Additionally, we could represent the domain as a collection of capsules, thus allowing for a more detailed domain geometry. We note that if we have more than three primitives, the domain might become disconnected, and it would be interesting to see how our bases could capture such scenarios.
Our current setup restricts the bases to align with the boundary, which is a typical computational fluid dynamics boundary condition. It would be interesting to extend our bases to support, for instance, no slip or an inflow/outflow boundary. This would require a user interface design to control the boundary condition and a mechanism to blend bases based on the input.

### v2

@@ 4.3. Ablation study
# Sample Points
Figure 9. Bases generated by omitting different parts of our pipeline. Each omission leads to different artifacts.
We showcase the importance of three main features of our pipeline (Figure 9).
Smoothness loss
This is the only loss that is not motivated by fundamental physical law. We included it to bias the MLP towards smooth solutions (which our bases must be). Adding it improves the reliability of the training process: in our experiments, without it, we obtain unreasonable bases more frequently. The last figure in the second row of Figure 9 shows that without the smoothness loss the MLP fails to generate a valid basis.
Rounded Corners
Number of Samples
We sample $10^{6}$ points for each input shape, as physics-based unsupervised losses require a sufficient number of samples to ensure convergence. Reducing the number of samples often leads to poor convergence, with the model violating physical constraints and producing trivial basis functions.
4.4. Animations
We run all our experiments on a GeForce RTX 3090. Generating the initial fluid is interactive, while the simulation runs at 40 frames per second for 250 thousand points. To visualize the flow, we render water displacement using the magnitude of the velocity field and show particle trajectories as they move along the flow. We refer to the additional materials for all the videos of our animations.
Two-dimensional Editing
@@ 4.4. Animations
$t=0.15s$
$t=0.25s$
$t=0.37$
Figure 10. Example of different fluid animation frames for different input sketches. The flow adapts to the position of the circles and the curves and produces natural animation. The top animation depicts a calm fluid surface, whereas the bottom demonstrates a turbulent scene with multiple vortices.
Moving Boundary
Our method naturally accommodates moving boundaries, as demonstrated by various examples involving translation and rotation for a single component and separate components, and even changes in the domain’s topology. These animations reveal how interesting fluid’s behaviors appears with such dynamics. Our method’s ability to handle moving boundaries makes it well-suited for integration into game engines. It can enrich user experience by enabling real-time generation of physically realistic fluid flows when characters interact with water, such as walking or moving through it.
Figure 1 bottom, where a duck moves around a stationary rock, shows an example of moving boundaries with multiple disconnected components using our time-dependent bases (Figure 11 top). From another angle (Figure 11 bottom) we show that, as the duck pushes into different corners, pre-existing vortices are compressed and eventually dissipated.
$t=0.025s$
$t=0.0375s$
$t=0.055s$
$t=0.07s$
Figure 11. A rubber duck moves around a stationary rock (bottom). As it pushes toward the bottom and left corners, the vortices in those regions are compressed and gradually vanish due to the interaction with the duck’s movement. The top shows the bases for frame 0.
We first illustrate the effect of a spinning fan at different speeds (Figure 12). When the fan spins rapidly, particles are pushed away from the blade region, effectively preventing them from entering. In contrast, with a slower rotation, particles are able to approach and circulate near the blades.
Slow fan.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fast fan.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$t=0s$
$t=0.25s$
$t=0.5s$
$t=0.75s$
$t=1s$
Figure 12. Two-dimensional fluid animation using our neural kinematic bases. We visualize the magnitude of the velocity as a height-field in the background, while we advect particles to better illustrate movement.
Figure 13 shows a gripper opening and closing, altering the genus of the domain. During the closing motion, both inward and outward flows appears and some particles are carried into the gripper while others are pushed out. Once the gripper is fully closed, the flow circulates around it, forming a loop and particles inside the gripper become trapped. When the gripper reopens, the flow pattern reverses, and the trapped particles are released while new particles are drawn in by the surrounding flow.
$t=0s$
$t=0.25s$
$t=0.5s$
$t=0.75s$
$t=1s$
Figure 13. Example of fluid animation where the domain changes genus. When the cavity appear, the flow becomes “trapped.” It is the released as it opens.
@@
All our methods can be generalized to three dimensions. The domain is now represented by ten spheres and their radii (40 parameters instead of 30); the MLP network accepts 3D points and produces a volumetric vector field (Figure 1 top shows some bases for a fan). Figure 1 middle shows an animation of a moving 3D fan where the flow moves around the blade, while Figure 14 shows an animation with static boundaries. Both figures show the particle traces from spherical sources (in different colors) advected by the velocity field.
@@
$t=9.65s$
Figure 14. Three-dimensional fluid animation using our neural kinematic bases of a fixed obstacle. We show particle traces from two spherical sources advected by the velocity field.
5. Conclusions
We introduce a novel neural kinematic base represented by an MLP that captures fundamental physical properties and generalizes to different domains. We show how using our bases and a standard advection integrator, we can generate two- and three-dimensional animations of moving obstacles.
To keep the training times reasonable, we decided to use only ten bases, but our method should be able to generate more bases. Additionally, we could represent the domain as a collection of capsules, thus allowing for a more detailed domain geometry.
Our method enforces Dirichlet boundary conditions, aligning the bases with the domain boundaries. Extending it to support other types, such as Neumann conditions, would require additional mechanisms to specify and blend boundary constraints. Our bases are inherently smooth, which produces smooth fluid animations and boundary geometries but limits applicability to non-smooth boundaries. Moreover, because Dirichlet conditions prevent prescribing velocity normal to obstacles, obstacles cannot actively push the fluid, resulting in quasi-static behavior. We believe that these phenomena are interesting and we leave their investigation as future work.
@@ 5. Conclusions
Acknowledgements.
The Bellairs Workshop on Computer Animation was instrumental in the conception of the research presented in this paper. This work was partially supported by the NSERC grants DGECR-2021-00461, RG-PIN 2021-03707, RGPIN-2024-04523, and RGPIN-2024-04605, FRQNT 365040, as well as a gift from Adobe Research. This work was also supported in part through the Digital Research Alliance of Canada resources, services, and staff expertise.
Appendix A Comparisons
Cui et al. (2018)
Ours
frame 0
frame 100
Figure 15. Comparison with a fixed fan as obstacle with velocity field initialized from a large smoke.
Cui et al. (2018)
Ours
frame 0
frame 100
Figure 16. Comparison with a fixed fan as obstacle with velocity field initialized from a small smoke.
We evaluate two 2D scenes with different smoke scales (figures 15 and 16), fitting the velocity field from Cui et al. (2018) to our bases. Since the two methods differ in their representational capacity, our approach performs less favorably when fitting fields with many zero values (small-scale case). However, it demonstrates better alignment with obstacle boundary conditions as our bases account for obstacle boundaries during generation rather than relying on post-processing.